OpenStudy (anonymous):

For which system of equations is (5, 3) the solution?

OpenStudy (anonymous):

\[x+y=8\\x-y=2\] would work

OpenStudy (anonymous):

A. 3x – 2y = 9 3x + 2y = 14 B. x – y = –2 4x – 3y = 11 C. –2x – y = –13 x + 2y = –11 D. 2x – y = 7 2x + 7y = 31

OpenStudy (anonymous):

how about any of these

OpenStudy (anonymous):

plug the numbers in and see which one works

OpenStudy (anonymous):

\[3x – 2y = 9\\ 3x + 2y = 14\] \[3\times 5-2\times 3=9\\ 3\times 5+2\times 3=14\] true or false?

OpenStudy (anonymous):

true?